OpenStudy (anonymous):

What are the factors of 3x2 + 13x + 4? (3x + 1)(x + 4) (3x + 4)(x + 1) (3x − 4)(x + 1) (3x − 1)(x + 4)

OpenStudy (anonymous):

3x2 + 13x + 4 =3x2 + 12x+x + 4 =3x(x+4) +1(x+4) =(x+4)(3x+1)

OpenStudy (anonymous):

= (3x + 1)(x + 4)

OpenStudy (anonymous):

Ty